Quantitative Problem: Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis. Both projects' after-tax cash flows are shown on the time line below. Depreciation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firm's average project. Bellinger's WACC is 9%.

0 1 2 3 4
Project A -1,030 640 340 260 310
Project B -1,030 240 275 410 760

What is the NPV of Project A and of Project B?
Answer #1

A:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=640/1.09+340/1.09^2+260/1.09^3+310/1.09^4

=1293.71

NPV=Present value of inflows-Present value of outflows

=1293.71-$1030

=$263.71(Approx).

B:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=240/1.09+275/1.09^2+410/1.09^3+760/1.09^4

=1306.64

NPV=Present value of inflows-Present value of outflows

=1306.64-1030

=$276.64(Approx).